An electron moves with a speed of 8.0 x 106 m/s along the +x-axis. It enters a region where there is a magnetic field of 2.5 T, directed at an angle of 60° to the +x-axis and lying in the r plane. (1 ev-1.60 x 10-19 C, mel = 9.11x 10-31 kg) Calculate the magnitude of (a) the magnetic force on the electron (b) the acceleration of the electron.
1
Expert's answer
2020-06-29T14:06:15-0400

a)


"F=eBv\sin{60}\\F=(1.6\cdot10^{-19})(2.5)(8\cdot10^{6})\sin{60}\\F=2.77\cdot10^{-12}N"

b)


"a=\frac{F}{m}=\frac{2.77\cdot10^{-12}}{9.11\cdot10^{-31}}=3.04\cdot10^{18}\frac{m}{s^2}"